Prev Next

SW&IT Meeting Minutes 2018-09-03

Attendees: Andrew, David, Igor, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

News from management

  • visitor instrument guidelines
  • s/w recruitment
  • UCAP

Previous Action Items

  • Vsphere (DE)
  • power plan update draft (DE)
  • LBTI/jumbo switch (DE) backup plan?
  • Luci H/O doc (IS)

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input form secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s) Primary Secondary Week
OAC
AC (2)
Petr (1)
1
AzCam
AC (1)
Petr 2
DIMM
AC (1)
Petr 2
Weatherstation AC Petr 3
FACSUM,WEATHERSTATION AC Matthieu 4
ALH Igor(1/2)
Matthieu (1)
1
INDI Igor Matthieu 3
GCS
Petr (2)
Matthieu (2)
1
Telemetry
Petr (1)
Matthieu 2
OVMS
Petr (2)
Matthieu (2)
3
Status Server Stephen Matthieu 2
Telemetry cleanup Stephen Petr 3
LBC
Yang (1)
Igor(1/2) 1
MODS Yang Igor 2
Telem-VIZ Yang Stephen 3
particle monitors Yang Petr 4

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • jumbo frame network
  • trivial passwords (ovms-uei, others?)
  • KS handovers
    • use Twiki vs. email
    • use Team Calendar for meeting schedule
  • KS other handovers, e.g.
  • Luci telemetry - web-io cold heads state (for vibration)
  • ovms startup default (JH email)
  • AGW4 g/w swap, w lenslet box coord issue

This week in github

Team reports

YZ

  • Working on transition to Kellee's LBC and MODS responsibility (focused on LBC currently).
  • Working with Petr on collimation and pointing analysis and improvements of procedures and programs.
  • TPoint software and old pointing script reorganization in Github.
  • Continue working on documenting pointing and collimation procedures.
  • Continue working on Python package and modules.
  • Need to start gather more info on Al's AO project.

AC

last week:
  • OAC and AGW overview and walkthrough
  • ARGOS and TAN build with tests

this week:
  • AzCam and DIMM knowledge transfer
  • FOSWiki transition
  • Revert MagExposure on AGW3
  • ARGOS handover edits

MB

  • management OC, WPC restarted
  • PP2019 meeting
  • Labor reporting roll out
  • KS Transition plan roll out
  • Science Archive xslt digging into PIT backups and OT logs
  • git migration tpoint, argos-tbad, argos-lch and others for PK
  • s/w interview continued
  • Hand Over prep GCS(2), ALH(1), OVMS(2)

next:
  • WPC chair
  • IT to github issues
  • AC/KS transition planning
  • science archive f ollow up
  • Hand Over prep ALH(2), INDI(1), Tele(1), StatsServ (1)

PK

  • GCS and OVMS transtion
  • prepared 2018B for instalation (binaries on mountain)
  • well spend collimation afternoon with Yang
  • tpoint, tpk, slalib and tcspk in GitHub (after OK from David and Patrick)

Weekend @ Lowell (Flagstaff):
  • DDW dome driver
  • crash course in GigE Vision, basic driver
  • Planeweve Elmo controllers

Next:
  • shedule GCS (and OVMS) presentations
  • 2018B installation and checkout (today)
  • AzCam and DIMM secondary
  • Skipping Telemetry discussions - I will go straight to TechTalk?

KS

Last week
  • Updated GCS on the mountain to be 2018A patched version for new HK. Tested both AGW2 and AGW4 (different types of HK). Had to update the OAC library on TCS machines. Took some time because TCS was also not in a good state after the power problems - gluster issues and mounts.
  • Modified the OAC build to not link the telemetry library into clients.
  • Set up rsync of OAC telemetry files to Tucson.
  • Updated notes for OVMS kernel update.
  • Checked in GCS updates for 2018B.
  • Released updated LBC version with just two IT updates.
  • Transition
    • OAC note to team for where to start, what it is. Short introductory discussion with AC since he and Dan had some issues.
    • LBC note to team, including the SPIE paper from 2014.
    • GCS/OVMS discussion with MB and PK.
  • Start up
    • AGW4 was going off - turns out the WFS camera threshold was low, even though Dan had suggested he was going to be very conservative and didn't want alarms in the beginning.
    • LBC looks good - an issue with testfilters in the beginning, but it may have been because nothing was up? I should look at it.
    • DIMM is good - they didn' have to modify the target for auto-acquisition.
    • OVMS has some bad channels, but it may be because they were disconnected for AGW work.

This week
  • Transition meetings - ALH, OAC, LBC.
  • Wiki updates for OAC telemetry and new build dependencies.
  • Find all my softcopy diagrams and put them somewhere.
CONFLICT version 18:

IS

This week
  • LUCI
    • LUCI RTD/Aladin development
    • Possibly AFC testing?
  • OT
    • PIT bug fixes
    • build binocular OT test version
  • ALH/INDI adoption
  • LBC/MODS?
    CONFLICT version new:

DE

Last week:
  • Set up meeting with VMWare to go over upgrade plan
  • Spoke to rep re: server room power. He is finding options in our search space for me
  • Jonathan noted buzzing noise from AGw computer power supply; I ordered spares
  • Set up SWG laptop for Petr's use
  • Un-mothballed several VMs from the archive at users' request
  • Discovered ordering error with DIN switches, have ordered the correct model and have CDW rep pushing for quick shipment
This week:
CONFLICT end
  • VMWare meeting Wednesday
  • Followup with Provantage re: UPS
  • Push DIN switch order
  • Swap AGw power supply
Topic revision: r22 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback